# Feigenbaum universality in subcritical Taylor-Couette flow

### v1

@@ Abstract
Feigenbaum universality is shown to occur in subcritical shear flows. Our testing ground is the counter-rotation regime of the Taylor-Couette flow, where numerical calculations are performed within a small periodic domain. The accurate computation of up to the seventh period doubling bifurcation, assisted by a purposely defined Poincaré section, has enabled the estimation of the two Feigenbaum universal constants with unprecedented accuracy in a fluid flow problem. We have further devised a method to predict the bifurcation diagram up to the accumulation point of the cascade based on the detailed inspection of just the first few period doubling bifurcations. Remarkably, the method is applicable beyond the accumulation point, with predictions remaining valid, in a statistical sense, for the chaotic dynamics that follows.
@@ 1 Introduction
The counter-rotating regime of Taylor-Couette flow has long served as a paradigm for subcritical transition to turbulence. The onset of fully developed turbulence is preceded by intermittency, the flow typically exhibiting alternating laminar and turbulent helical bands (Coles, 1965; Andereck et al., 1986; Meseguer et al., 2009b; Dong, 2009) whose mean flow topology (Wang et al., 2023) has recently been shown similar to the laminar-turbulent stripe pattern observed for plane Couette flow (Barkley & Tuckerman, 2007). The multi-scale nature of the stripe structure complicates the detailed understanding of subcritical transition with all the subtle features it encompasses. Nevertheless, the aspects associated with the emergence of chaos can often be approached in small periodic domains called minimal flow units (Kawahara et al., 2012). A recent review paper (Feldmann et al., 2023) hypothesised a connection between subcritical Taylor-Couette flow and plane Couette flow, for which Kreilos & Eckhardt (2012) reported a period-doubling cascade leading to chaos in a small periodic box. Here, we confirm that a period-doubling route to chaos does indeed occur in Taylor-Couette flow, within the parameter range studied by Meseguer et al. (2009b), Deguchi et al. (2014), and Wang et al. (2022), hereafter abbreviated as W22. The minimal periodic domain we use has an annular-parallelogram shape (Deguchi & Altmeyer, 2013) with two of its sides aligned along the helical band. As reported in W22, the use of this domain was key to observing the first few period-doubling bifurcations.
@@
Our preliminary computations suggested that the subsequent period doubling cascade may align exceptionally well with Feigenbaum universality (Feigenbaum, 1978, 1979, 1980, 1982), thus inspiring the detailed exploration we perform here. This type of universality became widely recognised among fluid dynamicists following the natural convection experiment conducted by Libchaber et al. (1982), which led to the first estimation of Feigenbaum’s first constant, $\delta_{\rm F}\approx 4.6692016$, corresponding to the limiting ratio of a period-doubling bifurcation interval to the next (see also the recent article by Libchaber, 2023). Their value, $\delta_{4}=4.4\pm 0.1$, estimated from up to the fourth period-doubling bifurcation point (hence the subindex $n=4$), was not too distant from the theoretical value $\delta_{\rm F}$. Buzug et al. (1993) found an estimate close to $\delta_{\rm F}$ when studying a Taylor-Couette apparatus with a tilted lid, while subsequent experimental efforts, summarised in table 1, resulted in persistently larger discrepancies.
@@
Rayleigh-Bénard (water) E $4$ $4.3$$\pm$ $0.8$ Giglio et al. (1981)
Rayleigh-Bénard (mercury) E $4$ $4.4$$\pm$ $0.1$ Libchaber et al. (1982)
@@
Table 1: Feigenbaum universality analyses in fluid systems. The table includes the number of period doubling bifurcations analysed ($n$), and the experimental (E) or numerical (N) nature of the study. Feigenbaum’s first constant, as estimated from the last three period doubling bifurcations analysed in each case is given in column $\delta_{n}$.
@@
Estimates of Feigenbaum’s first universal constant have also been obtained numerically for various fluid systems. Some of the values reported for $\delta_{n}$, also listed in table 1, are reasonably close to $\delta_{\rm F}$ at first glance. However, it must be borne in mind that the reliability of the estimates depends strongly on the number of consecutive period-doubling bifurcations analysed and is extremely sensitive to the accuracy with which they are located. Unfortunately, the estimates of $\delta_{\rm F}$ we have found in the literature of fluid flows were not accompanied with a detailed analysis as done, for example, for the Kuramoto-Sivashinsky equation (Smyrlis & Papageorgiou, 1991). Accordingly, their reliability is debatable, as concluding Feigenbaum universality from a handful of period-doubled solutions is, at the very least, perilous.
@@
All the studies presented in table 1 explore parameter ranges for which the respective base flows are unstable, while our focus is set on the subcritical regime. The aforementioned study by Kreilos & Eckhardt (2012) was the first to show that period-doubling cascades may also occur in subcritical transition problems. Their approach consisted in selecting the smallest possible periodic domain that is capable of sustaining turbulence in plane Couette flow while, at the same time, imposing specific discrete symmetries to further constrain the dynamics. Symmetry restrictions were also employed by Moore et al. (1983) and Van Veen (2005) in their respective works. A similar period-doubling cascade in plane Couette flow has also been reported in a more recent paper by Lustro et al. (2019). However, calculation of $\delta_{\rm F}$ in the subcritical regime of fluid flow problems has not been conducted to date.
@@
Although period-doubling cascades have been known to occur for over a century (Collet, 2019), universality was not discovered until much later by Feigenbaum (1978) and Coullet & Tresser (1978) independently, hence its being often referred to as Feigenbaum-Coullet-Tresser universality by mathematicians. The existence of universality swiftly disseminated throughout the mathematical community, as vividly depicted by Khanin et al. (2021). Since its unveiling, universality has been repeatedly observed in numerical simulations of low-dimensional systems of ordinary differential equations. Checking universality in low-dimensional models like the Rössler system or the Duffing equation has become a common academic exercise. Feigenbaum universality was originally established within the framework of discrete dynamical systems based on one-dimensional iterated maps of the form $x_{\ell+1}=f(x_{\ell})$, $\ell\in\mathbb{N}$, with $f(x)$ a unimodal function. Its necessary occurrence in continuous time dynamical systems is typically justified by the use of Poincaré sections. The phenomenon is often illustrated by initially replacing the Smale horseshoe that occurs on the Poincaré section with a Hénon map, which reduces to the logistic map in the limiting case of vanishing area contraction rate after one mapping. Then, since the logistic map is a unimodal map, universality of the period doubling cascade can be explained by means of renormalisation theory, as done in many standard textbooks such as Collet & Eckmann (1980); Schuster (1988); Glendinning (1994) or Strogatz (2024). Universality manifests in both the parameter (coordinate) and state (ordinate) axes of the bifurcation diagram, the scaling in the latter axis bearing close relation with Feigenbaum’s second constant, $\alpha_{\rm F}\approx-2.5029079$. In renormalisation theory, $\alpha_{\rm F}$ and the Feigenbaum function $G$ uniquely solve the so-called Feigenbaum-Cvitanović functional equation $\mathcal{R}[G]=G$ (Feigenbaum, 1978; Cvitanović, 1989), with the condition $G(0)=1$. Here, the action of the renormalisation operator, $\mathcal{R}$, on a function $f$ consists merely in the twice repeated application of $f$, mediated by a rescaling that involves a factor $\alpha_{\rm F}$:
@@
Despite its simplicity, this operator lies at the heart of the various instances of self-similarity that arise in the bifurcation diagram of the map generated by $f$. Likewise, Feigenbaum’s first constant $\delta_{\rm F}$ is the leading eigenvalue of the linearisation around $G$ of the renormalisation operator $\mathcal{R}$, and $\varPhi$ is the associated eigenfunction (Lanford III, 1982; Epstein, 1986; Eckmann & Wittwer, 1987; Briggs, 1991; Sthephenson & Wang, 1991), which has recently been dubbed Feigenfunction by Thurlby (2021), as a tribute to Feigenbaum. The first mathematical proof of the existence of the universal constants, $\delta_{\rm F}$ and $\alpha_{\rm F}$, and functions, $G$ and $\varPhi$, is attributed to Lanford III (1982).
@@
Ascertaining universality from the estimation of $\delta_{\rm F}$ and $\alpha_{\rm F}$ in fluid flow problems presents a twofold challenge. First, a numerical analysis analogous to that of low-dimensional models demands significantly higher computational resources when dealing with the Navier-Stokes equations. The first estimation of $\delta_{\rm F}$ in the field of computational fluid mechanics was contributed by Moore et al. (1983), when analysing two-dimensional thermosolutal convection. It took, however, until Van Veen (2005) to expose universality in a three dimensional setup, and even then, only for a fluid contained in the simplest computational domain, i.e., periodic in all three space directions. Second, the governing equations are very high-dimensional (formally infinite dimensional before spatial discretisation) and thus the link to one-dimensional maps is far less obvious than for low-dimensional dynamical systems. Feigenbaum universality does not apply unconditionally to high-dimensional systems exhibiting period-doubling cascades. Cascades sequentially doubling the period at every bifurcation may occur through mechanisms quite different from those leading to universality (e.g., Yalim et al., 2019). Other routes to chaos are also common in fluid systems; in the Taylor-Couette system, for example, chaos has been found to arise following the Ruelle-Takens-Newhouse scenario (Swinney & Gollub, 1985), or Shil’nikov-type bifurcations (Lopez & Marques, 2005). Therefore, detecting the onset of chaos, which requires a great deal of effort, does not necessarily indicate that a universal period-doubling cascade has taken place. All these difficulties justify why demonstrating Feigenbaum universality in a fluid system requires careful analysis of the phase space. Buzug et al. (1993) reconstructed the dynamics in phase space via delay-time embedding, while Kreilos & Eckhardt (2012) and Gao et al. (2015) investigated the relationship between consecutive local maxima of the time series of some physical quantity. Even though both ways of sampling a continuous system to produce a discrete system are valid, the use of a purposely defined Poincaré section provides the most reliable approach.
@@
The stable orbits of ever increasing period that arise in succession as the parameter is varied along the period-doubling cascade accumulate at the so-called accumulation point, beyond which chaotic dynamics ensue. Our interest extends also to phenomena occurring past this point. The existence of a reverse cascade, whereby chaotic bands successively merge in pairs as one moves away from the accumulation point, is a well-established property of simple model maps (Grossmann & Thomae, 1977; Lorenz, 1980). It is noteworthy that Huberman & Rudnick (1980) observed self-similarity in the numerical analysis of Lyapunov exponents at the onset of chaos. Nevertheless, the extent to which universality holds in the chaotic regime remains an elusive question. Libchaber (1983) tried to detect universality beyond the accumulation point in a natural convection experiment. To the authors’ knowledge, this is the first and only endeavour to unveil universality past the accumulation point of a period-doubling cascade in a fluid dynamics problem, but the results were rather crude due to experimental limitations. In theory, if universality holds beyond the accumulation point, the bifurcation diagram should be predictable, not only up to but also past this point, from the initial few period-doubling bifurcations. However, we have not been able to find any such attempts in the literature, even for low-dimensional models.
@@
The paper is structured as follows. The problem formulation is presented in §2, alongside a brief account of the numerical methods employed. §3 introduces the period doubling cascade that is our object of study and illustrates the procedure by which period doubling bifurcation points are accurately computed. The sequence of the first few such points is then exploited in §4 to estimate the first and second Feigenbaum constants and to extrapolate the expected occurrence of the accumulation point. A method for the detailed prediction of the bifurcation diagram in the neighbourhood of the accumulation point from just a few initial period doubling bifurcations is devised in §5. We further show how meaningful predictions can be made, in a statistical sense, that extend beyond the accumulation point into the chaotic regime. Finally, the main findings are summarised and conclusions drawn in §6.

### v2

@@ Abstract
Feigenbaum universality is shown to occur in subcritical shear flows. Our testing ground is the counter-rotation regime of the Taylor-Couette flow, where numerical calculations are performed within a small periodic domain. The accurate computation of up to the seventh period doubling bifurcation, assisted by a purposely defined Poincaré section, has enabled us to reproduce the two Feigenbaum universal constants with unprecedented accuracy in a fluid flow problem. We have further devised a method to predict the bifurcation diagram up to the accumulation point of the cascade based on the detailed inspection of just the first few period doubling bifurcations. Remarkably, the method is applicable beyond the accumulation point, with predictions remaining valid, in a statistical sense, for the chaotic dynamics that follows.
@@ 1 Introduction
The Taylor-Couette flow, driven by two independently rotating coaxial cylinders (figure 1a), has been a prominent research topic in fluid dynamics for over a century. By varying the angular velocities of the cylinders in experimental setups, a rich variety of flow patterns can be observed, as documented by Andereck et al. (1986). In this paper, we focus on the case with counter-rotating cylinders, a regime for which Taylor-Couette flow has long served as a paradigm of subcritical transition to turbulence. The onset of fully developed turbulence is preceded by intermittency, with the flow typically exhibiting alternating laminar and turbulent helical bands (Coles, 1965; Meseguer et al., 2009b; Dong, 2009); see figure 1b.
(a) (b)
Refer to caption Refer to caption
Figure 1: Taylor-Couette flow. (a) Sketch of the flow configuration. The inner and outer cylinders have radii $r=r_{i}$ and $r=r_{o}$, respectively, and rotate with angular velocities $\Omega_{i}$ and $\Omega_{o}$. (b) A snapshot of the stripe pattern adopted from figure 2 of Wang et al. (2022). The colour map shows the radial vorticity at the mid gap $r_{m}=(r_{o}+r_{i})/2$. The radius ratio and inner and outer cylinder Reynolds numbers, defined in §2, are set to $(\eta,R_{i},R_{o})=(0.883,600,-1200)$.
The mean flow topology of these helical bands (Wang et al., 2023) has recently been shown to resemble the laminar-turbulent stripe pattern observed in plane Couette flow (Barkley & Tuckerman, 2007). Indeed, large-scale intermittent patterns of coexisting turbulent and laminar regions are ubiquitous in wall-bounded subcritical shear flows (Tuckerman et al., 2020). Stochastic approaches, such as directed percolation theory, are gaining popularity as a framework for understanding the onset of subcritical turbulent transition (Hof, 2023). However, applying stochastic arguments to a deterministic system inherently assumes the pre-existence of chaotic dynamics. In the case of subcritical shear flows, identifying the emergence of chaos that precedes turbulent transition, poses a significant challenge.
@@
What complicates the detailed understanding of transitional and fully developed turbulence in shear flows, with all the subtle features it encompasses, is the multi-scale nature of the flow. Consequently, many aspects of subcritical parallel shear flow turbulence have historically been approached using small periodic computational domains called minimal flow units (Jiménez & Moin, 1991; Hamilton et al., 1995; Kawahara et al., 2012). In Taylor-Couette flow, the use of minimal boxes has been common since the early stages of simulations (e.g. Marcus (1984); Coughlin & Marcus (1992)). However, most studies employed parameters in the supercritical regime, and somewhat surprisingly, research in the subcritical regime remains rare. A recent review paper (Feldmann et al., 2023) hypothesised a connection between subcritical Taylor-Couette flow and plane Couette flow, for which Kreilos & Eckhardt (2012) reported a period-doubling cascade leading to chaos in a small periodic box. Here, we confirm that a period-doubling route to chaos does indeed occur in Taylor-Couette flow within the parameter range studied by Meseguer et al. (2009b), Deguchi et al. (2014), and Wang et al. (2022), the latter hereafter abbreviated as W22.
The most suitable flow unit for the parameter regime we address here is of annular-parallelogram shape proposed by Deguchi & Altmeyer (2013). W22 first used a narrow elongated domain (red box in figure 1b) to compute turbulent stripes in their minimal natural box, a strategy that has been repeatedly adopted in plane Couette (Barkley & Tuckerman, 2007) and channel flows (Tuckerman et al., 2014). The domain was subsequently shortened in the azimuthal direction (cyan box in figure 1b) so as to fit only one streamwise wavelength of a typical coherent structure of developed turbulence. As reported by W22, the use of such a domain was instrumental in observing the first few period-doubling bifurcations in the route to chaos. In this paper, we show that the ensuing period doubling cascade aligns exceptionally well with Feigenbaum universality (Feigenbaum, 1978, 1979, 1980, 1982).
1.1 Feigenbaum cascade in fluid flows
@@
Feigenbaum universality became widely recognised among fluid dynamicists following the natural convection experiment conducted by Libchaber et al. (1982), which led to the first empirical approximation of Feigenbaum’s first constant, $\delta_{\rm F}\approx 4.6692016$, corresponding to the limiting ratio of a period-doubling bifurcation interval to the next (see also the recent article by Libchaber, 2023). Their value, $\delta_{4}=4.4\pm 0.1$, estimated from up to the fourth period-doubling bifurcation point (hence the subscript $n=4$), was not too distant from the theoretical value $\delta_{\rm F}$. Buzug et al. (1993) found an estimate close to $\delta_{\rm F}$ when studying a Taylor-Couette apparatus with a tilted lid, while subsequent experimental efforts, summarised in table 1, resulted in persistently larger discrepancies.
@@
Thermal convection E $4$ $4.3$$\pm$ $0.8$ Giglio et al. (1981)
Thermal convection E $4$ $4.4$$\pm$ $0.1$ Libchaber et al. (1982)
@@
Table 1: Feigenbaum universality analyses in fluid systems. The table includes the number of period doubling bifurcations analysed ($n$), and the experimental (E) or numerical (N) nature of the study. An approximation to Feigenbaum’s first constant, estimated from the last three period doubling bifurcations analysed in each case, is given in column $\delta_{n}$.
@@
Estimates compatible with Feigenbaum’s first universal constant have also been obtained numerically for various fluid systems. Some of the values reported for $\delta_{n}$, also listed in table 1, are reasonably close to $\delta_{\rm F}$ at first glance. However, it must be borne in mind that the reliability of the approximations depends strongly on the number of consecutive period-doubling bifurcations analysed and is extremely sensitive to the accuracy with which they are located. Unfortunately, the approximate determinations of $\delta_{\rm F}$ we have found in the literature of fluid flows were not accompanied with a detailed analysis as done, for example, for the Kuramoto-Sivashinsky equation (Smyrlis & Papageorgiou, 1991). Accordingly, their reliability is debatable, as concluding Feigenbaum universality from a handful of period-doubled solutions is, at the very least, perilous.
@@
All the studies presented in table 1 explore period doubling cascades that follow from a supercritical sequence of bifurcations of the base laminar flow. We focus instead on the subcritical regime, where turbulent transition may occur despite the linear stability of the base flow and can only be triggered by finite amplitude perturbations. The detection and study of period doubling bifurcations emanating from unstable solutions is impracticable, and stable finite-amplitude solutions in the subcritical regime of shear flows that could potentially originate a cascade are rare and hard to find. The aforementioned study by Kreilos & Eckhardt (2012) was the first to show that period-doubling cascades may also occur in subcritical transition problems. Their approach consisted in selecting the smallest possible periodic domain that is capable of sustaining turbulence in plane Couette flow while, at the same time, imposing specific discrete symmetries to further constrain the dynamics. Unfortunately, the resolution of their parametric exploration was insufficient to positively confirm universality. Symmetry restrictions were also employed by Moore et al. (1983) and Van Veen (2005) in their respective works. A similar period-doubling cascade in plane Couette flow has also been reported in a more recent paper by Lustro et al. (2019). However, calculation of $\delta_{\rm F}$ in the subcritical regime of fluid flow problems has not been conducted to date.
1.2 Mathematical aspects of the Feigenbaum cascade
@@
Although period-doubling cascades have been known to occur for over a century (Collet, 2019), universality was not discovered until much later by Feigenbaum (1978) and Coullet & Tresser (1978) independently, hence its being often referred to as Feigenbaum-Coullet-Tresser universality by mathematicians. The existence of universality swiftly spread throughout the mathematical community, as vividly depicted by Khanin et al. (2021). Since its unveiling, universality has been repeatedly observed in numerical simulations of low-dimensional systems of ordinary differential equations. Checking universality in low-dimensional models like the Rössler system or the Duffing equation has become a common academic exercise. Feigenbaum universality was originally established within the framework of discrete dynamical systems based on one-dimensional iterated maps of the form $x_{\ell+1}=f(x_{\ell})$, $\ell\in\mathbb{N}$, with $f(x)$ a unimodal function. Its necessary occurrence in continuous time dynamical systems is typically justified by the use of Poincaré sections. The phenomenon is often illustrated by initially replacing the Smale horseshoe that occurs on the Poincaré section with a Hénon map, which reduces to the logistic map in the limiting case of vanishing area contraction rate after one mapping. Then, since the logistic map is a unimodal map, universality of the period doubling cascade can be explained by means of renormalisation theory, as done in many standard textbooks such as Collet & Eckmann (1980); Schuster (1988); Glendinning (1994) or Strogatz (2024). Universality shows in both the parameter (coordinate) and state (ordinate) axes of the bifurcation diagram, the scaling in the latter axis is related to Feigenbaum’s second constant, $\alpha_{\rm F}\approx-2.5029079$. In renormalisation theory, $\alpha_{\rm F}$ and the Feigenbaum function $G$ uniquely solve the Feigenbaum-Cvitanović functional equation $\mathcal{R}[G]=G$ (Feigenbaum, 1978; Cvitanović, 1989), with the condition $G(0)=1$. Here, the action of the renormalisation operator, $\mathcal{R}$, on a function $f$ consists merely in the twice repeated application of $f$, mediated by a rescaling that involves a factor $\alpha_{\rm F}$:
@@
Despite its simplicity, this operator lies at the heart of the various instances of self-similarity that arise in the bifurcation diagram of the map generated by $f$. Likewise, Feigenbaum’s first constant $\delta_{\rm F}$ is the leading eigenvalue of the linearisation around $G$ of the renormalisation operator $\mathcal{R}$, and $\varPhi$ is the associated eigenfunction (Lanford III, 1982; Epstein, 1986; Eckmann & Wittwer, 1987; Briggs, 1991; Stephenson & Wang, 1991), which has recently been dubbed Feigenfunction by Thurlby (2021), as a tribute to Feigenbaum. The first mathematical proof of the existence of the universal constants, $\delta_{\rm F}$ and $\alpha_{\rm F}$, and functions, $G$ and $\varPhi$, is attributed to Lanford III (1982).
@@
Ascertaining universality from the ability to reproduce $\delta_{\rm F}$ and $\alpha_{\rm F}$ in fluid flow problems presents a twofold challenge. First, a numerical analysis analogous to that of low-dimensional models demands significantly higher computational resources when dealing with the Navier-Stokes equations. The first observation of $\delta_{\rm F}$ in the field of computational fluid mechanics was contributed by Moore et al. (1983), when analysing two-dimensional thermosolutal convection. It took, however, until Van Veen (2005) to expose universality in a three dimensional setup, and even then, only for a fluid contained in the simplest computational domain, i.e., periodic in all three space directions. Second, identifying flow configurations that exhibit the specific route to chaos being targeted, among the several possible in multi-dimensional systems, is much more difficult in physically relevant problems than in engineered toy models with tunable parameters. Cascades sequentially doubling the period at every bifurcation may occur through mechanisms quite different from those leading to universality (e.g., Yanagida, 1987; Kokubu et al., 1996; Homburg et al., 2001; Yalim et al., 2019). Other routes to chaos are also common in fluid systems; in the Taylor-Couette system, for example, chaos has been found to arise following the Ruelle-Takens-Newhouse scenario (Swinney & Gollub, 1985) , or Shil’nikov-type bifurcations (Lopez & Marques, 2005). All these difficulties justify why demonstrating Feigenbaum universality in a fluid system requires careful analysis.
1.3 The accumulation point of the cascade and beyond
@@
The orbits of ever increasing period that arise in succession as the parameter is varied along the period-doubling cascade pile up at the accumulation point, beyond which chaotic dynamics ensue. Our interest extends also to phenomena occurring past this point. The existence of a reverse cascade, whereby chaotic bands successively merge in pairs as one moves away from the accumulation point, is a well-established property of simple model maps (Grossmann & Thomae, 1977; Lorenz, 1980). It is noteworthy that Huberman & Rudnick (1980) observed self-similarity in the numerical analysis of Lyapunov exponents at the onset of chaos. Nevertheless, the extent to which universality holds in the chaotic regime remains an elusive question. Libchaber (1983) tried to detect universality beyond the accumulation point in a natural convection experiment. To the authors’ knowledge, this is the first and only endeavour to unveil universality past the accumulation point of a period-doubling cascade in a fluid dynamics problem, but the results were rather crude due to experimental limitations. In theory, if universality holds beyond the accumulation point, the bifurcation diagram should be predictable, not only up to but also past this point, from the initial few period-doubling bifurcations. However, we have not been able to find any such attempts in the literature, even for low-dimensional models. While self-similarity of the reverse cascade past the accumulation point has been invariably observed and has been conjectured to hold universally, the underlying mechanisms remain unknown.
1.4 Outline of the paper
@@
The paper is structured as follows. The problem formulation is presented in §2, alongside a brief account of the numerical methods employed. §3 introduces the period doubling cascade that is our object of study and illustrates the procedure by which period doubling bifurcation points are accurately computed. The sequence of the first few such points is then exploited in §4 to assess agreement with the first and second Feigenbaum constants and to extrapolate the expected occurrence of the accumulation point. A method for the detailed prediction of the bifurcation diagram in the neighbourhood of the accumulation point from just a few initial period doubling bifurcations is devised in §5. We further show how meaningful predictions can be made, in a statistical sense, that extend beyond the accumulation point into the chaotic regime. Finally, the main findings are summarised and conclusions drawn in §6.